# Computing with D-Algebraic Sequences

### v1

@@ Abstract
A sequence is difference algebraic (or D-algebraic) if finitely many shifts of its general term satisfy a polynomial relationship; that is, they are the coordinates of a generic point on an affine hypersurface. The corresponding equations are called algebraic difference equations (ADE). We show that subsequences of D-algebraic sequences, indexed by arithmetic progressions, satisfy ADEs of the same orders as the original sequences. Additionally, we provide algorithms for operations with D-algebraic sequences and discuss the difference-algebraic nature of holonomic and $C^{2}$-finite sequences.
@@
Affine D-algebraic sequence, subsequence, radical-rational dynamical system
@@ 1 Introduction
to accelerate the rate of convergence to $\lim_{n\to\infty}s(n)<\infty$. As discussed in [Wen89], this transformation has numerous applications in theoretical physics and other sciences [BJ75, Bre85]. This paper presents a general study of a class of sequences that remains unchanged by such transformations. These sequences satisfy nonlinear rational or polynomial recursions, similar to the one presented in (1).
@@
As an example, consider approximating $\sqrt{3}$ using the Babylonian method. The iteration is governed by the recursion
$s(n+1)=\frac{1}{2}\left(s(n)+\frac{3}{s(n)}\right),$ (2)
@@
with the chosen initial term $s(0)=1$. Using our result, we can systematically show that the Aitken transformation yields the sequence with initial term $t(0)=\frac{9}{5}$ and recursion
$t(n+1)=\frac{6\,t(n)}{t(n)^{2}+3}=\frac{6}{t(n)+\frac{3}{t(n)}}.$ (3)
The above may be generalized to any square root computation. Equations similar to (2) also appear in modeling, as illustrated with the discrete May-Leonard model by Roeger and Allen [RA04].
@@
Following J. F. Ritt’s algebraic approach to differential equations [Rit50], the study of associated objects such as differential polynomials and differential varieties was further developed by Raudenbush [Rau34], Rosenfeld [Ros59], and, more comprehensively with modern mathematics, by Kolchin [Kol73]. For difference equations, which relate to the classical $\Delta$ operator defined as $\Delta f(x)\coloneqq f(x+1)-f(x)$, R. Cohn developed much of the theory [Coh65].
@@
A key result of R. Cohn states that every nontrivial ordinary algebraically irreducible difference polynomial admits an abstract solution [Coh48]. This theorem has important implications for sequences over an algebraically closed field, say $\mathbb{K}$, of characteristic zero. Indeed, specialized to the difference ring $\mathbb{K}^{\mathbb{N}}$ (with the monoid of natural integers =$\mathbb{N}=\{0,1,\ldots\}$), Cohn’s theorem ensures the existence of sequences that are zeros of ordinary difference polynomial. We call them difference-algebraic (or D-algebraic) sequences and denote their set by $\mathbb{K}^{\aleph_{0}}$ (see Definition 2.8).
@@
D-algebraic functions [AEMSTT24, TT25] generalize functions defined by polynomial differential equations, encompassing classes such as D-finite functions [Kau23] and differentially definable functions [JPPS20] (see also [TTK21] for related power series). Similarly, D-algebraic sequences generalize sequences defined by polynomial equations involving their shifts and the index variable. These sequences arise naturally in computer science, for example, in the context of cost-register automata [ADD+13] and polynomial automata [BDSW17] over a unary alphabet. A particularly important subclass of D-algebraic sequences consists of the simple-rational recursive sequences, which are the zeros of difference polynomials linear in their highest shifts (l.h.s. difference equations) [CDBMP23]. It was shown in [TTW24] that any D-finite (holonomic or P-recursive) sequence satisfies an l.h.s. difference equation of order bounded by the sum of its holonomic degree and order. We provide a corrected proof of this statement here, as an error was inadvertently introduced in the final version of [TTW24]. Another notable subclass of D-algebraic sequences is that of $C^{2}$-finite sequences, introduced by Jiminez-Pastor, Nuspl, and Pillwein in [JPNP23]. We will give a constructive proof of their D-algebraicity.
This paper demonstrates that difference-algebraic sequences behave well under field operations (with a mild assumption for division) thanks to our algorithm for performing arithmetic operations on them. Our method is based on the differential elimination-prolongation method developed by Ovchinnikov, Pogudin, and Vo [OPV16], which was adapted to the difference algebra setting in [OPS20] (see also [GVDHYZ09]). Wibmer introduced a notion of dimension for systems of algebraic difference equations with solutions in sequence rings, showing that this dimension need not be an integer [Wib21]. We emphasize that the consistency of the systems of algebraic difference equations resulting from our arithmetic operations follows from Cohn’s existence theorem applied to the individual input difference polynomials, thereby avoiding the undecidability results of [PSW20]. While the computational theory for the differential case is relatively well-established (see, e.g., [BLOP95, Rob14, vDH19, TT25] and references therein), an effective theory for the arithmetic of (nonlinear) difference-algebraic sequences has received less attention. This motivates the present work. In addition to field operations, we prove that D-algebraic sequences are closed under partial sums, partial products, radicals, and composition with arithmetic progressions. This last result can be extended to composition with broader classes of strictly increasing integer sequences. We are unaware of comparable results for such compositions, and we consider our method for computing equations for D-algebraic subsequences a key contribution.
We conclude this introduction with a simple example illustrating our algorithm for the arithmetic of D-algebraic sequences.
Example 1.1.
We here compute an algebraic difference equation satisfied by the product of a Fibonacci sequence, i.e., a solution of the equation $s(n+2)=s(n+1)+s(n)$, and $k^{2^{n}}$, for some constant $k$, i.e., a solution of the equation $s(n+1)=s(n)^{2}$. Using our algorithm, we find the following third-order algebraic difference equation {dmath} -s(n+3)^3 s(n)^6-s(n+3)^2 s(n+2)^2 s(n+1) s(n)^4+s(n+2)^7 s(n)^2 -3 s(n+3) s(n+2)^4 s(n+1)^2 s(n)^2-5 s(n+3)^2 s(n+2) s(n+1)^4 s(n)^2 +s(n+3)^2 s(n+1)^7-4 s(n+3) s(n+2)^3 s(n+1)^5+4 s(n+2)^6 s(n+1)^3=0.
The above equation provides three possible continuations for the sequence, but the original two equations determine the correct one, thus defining a recursion for the entire sequence. The results in [CDBMP23] imply that whenever we have two initial equations that are linear in their highest shifts (l.h.s.), a higher-order l.h.s. difference equation can be derived for the arithmetic of their zeros. In this case, we obtain a fourth-order l.h.s. equation. $\blacksquare$
@@ 2 Definitions
We primarily consider the $\sigma$-ring $(\mathbb{K},\sigma)$, where $\mathbb{K}$ is an algebraically closed field of characteristic zero and $\sigma$ is the identity on $\mathbb{K}$.
@@
The ring of difference polynomials in one difference indeterminate $x$ over $\mathbb{K}$ , also called the free difference $\mathbb{K}$-algebra in $x$, is the difference ring $\left(\mathbb{K}[x_{0},x_{1},x_{2},\ldots],\tilde{\sigma}\right)$, where $\tilde{\sigma}$ extends $\sigma$ as follows: $\tilde{\sigma}(a)=\sigma(a)=a$, for any $a\in\mathbb{K}$, and $\tilde{\sigma}(x_{j})=x_{j+1}$, $j\in\mathbb{N}$. For simplicity, this difference ring is denoted by $\mathcal{D}_{\sigma}(\mathbb{K},x)\coloneqq\mathbb{K}\left[\sigma^{j}(x)\mid j\in\mathbb{N}\right]$ or $\mathbb{K}\left[\sigma^{\infty}(x)\right]$ with the property $\sigma^{j+1}(x)=\sigma(\sigma^{j}(x)),j\in\mathbb{N}$.
@@
A difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ is said to be in normal form when $x$ effectively occurs as the smallest shift in $p$. The order of a difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, denoted $\operatorname{ord}(p)$, is the maximum $j\in\mathbb{N}$ such that $\sigma^{j}(x)$ appears in the normal form of $p$. The degree of $p$ denoted $\deg(p)$ is the total degree of $p$ (in normal form) as a polynomial in the algebraic polynomial ring $\mathbb{K}[x,\sigma(x),\ldots,\sigma^{\operatorname{ord}(p)}(x)]$. We say that $p$ is l.h.s. (linear in its highest shift) when the degree of $p$ with respect to $\sigma^{\operatorname{ord}(p)}(x)$ is $1$.
Observe that the order of a difference polynomial is unchanged by application of $\sigma$. To any difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, we associate a recurrence equation called algebraic difference equation obtained by the equality $p=0$.
@@
We give the orders and degrees of the input recurrence equations in Example 1.1. We can always assume that $\mathbb{K}$ is the field of complex number $\mathbb{C}$. For $p\coloneqq\sigma^{2}(x)-\sigma(x)-x,\operatorname{ord}(p)=2$ and $\deg(p)=1$. For $q\coloneqq\sigma^{2}(x)-\sigma(x)^{2}$, $\operatorname{ord}(q)=1$ and $\deg(q)=2$. $\blacksquare$
In classical algebraic geometry, zeros of polynomials over $\mathbb{K}$ correspond to points in $\mathbb{K}^{N}$, the affine $N$-space over $\mathbb{K}$, for some fixed $N\in\mathbb{N}\setminus\{0\}$, defining algebraic sets or varieties with the ideals of those polynomials. In our setting, the corresponding points may be regarded with denumerable coordinates as we want any zero of $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ to have coordinates that vanish $\sigma^{j}(p)$, for all $j\in\mathbb{N}$. Every such point defines a sequence over $\mathbb{K}$. We denote the set of such sequences over $\mathbb{K}$ by $\mathbb{K}^{\aleph_{0}}$, where $\aleph_{0}$ is aleph zero, the cardinality of $\mathbb{N}$. This makes sense of the passage from a finite to an infinite (countable) number of coordinates. Note that $\mathbb{K}^{\aleph_{0}}\subset\mathbb{K}^{\mathbb{N}}$, and the inclusion is strict since $(n^{n})_{n\in\mathbb{N}}$ is not a zero of a difference polynomial. In Section 3.1, we will see that $\mathbb{K}^{\aleph_{0}}$ has the structure of a difference ring. This fact is used in Definition 2.6. This means that ring operations are defined componentwise, and for $(s(n))_{n\in\mathbb{N}}=(s(0),s(1),s(2),\ldots)\in\mathbb{K}^{\aleph_{0}}$ (we also use the notation $(s(n))_{n}$), we consider the endomorphism $\sigma^{\prime}$ such that $\sigma^{\prime}((s(n))_{n\in\mathbb{N}})=(s(n+1))_{n\in\mathbb{N}}$. This $\sigma^{\prime}$ is often known as the shift operator. Observe that $\Delta=\sigma^{\prime}-\text{id}$. The following definition helps to clarify the correspondence between the zeros of difference polynomials and $\mathbb{K}^{\aleph_{0}}$.
@@
Let $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ and $(s(n))_{n\in\mathbb{N}}\in(\mathbb{K}^{\aleph_{0}},\sigma^{\prime})$ as previously defined. We say that $(s(n))_{n\in\mathbb{N}}$ is a generic zero of $p$ if, under the unique homomorphism of difference rings $\mathcal{D}_{\sigma}(\mathbb{K},x)\longrightarrow~{}(\mathbb{K}^{\aleph_{0}},\sigma^{\prime})$ given by the extension that sends $x$ to $(s(n))_{n\in\mathbb{N}}$, $p$ is sent to 00.
We emphasize that our choice of the ground difference field $\mathbb{K}$ with the identity endomorphism puts no distinction between the concept of solutions, formal solutions, or generic zeros of difference polynomials over $\mathbb{K}$; see the details in [Lev08, Section 2.2].
Finally, we can define a difference-algebraic sequence based on the following corollary of R. Cohn’s existence theorem for ordinary difference polynomials.
@@
[Coh48, Theorem IV] Every difference polynomial in $\mathcal{D}_{\sigma}(\mathbb{K},x)\setminus\mathbb{K}$ has a generic zero.

### v2

@@ Abstract
A sequence is difference algebraic (or D-algebraic) if finitely many shifts of its general term satisfy a polynomial relationship; that is, they are the coordinates of a generic point on an affine hypersurface. The corresponding equations are denoted algebraic difference equations (ADEs). We propose a formal definition of D-algebraicity for sequences and investigate algorithms for their closure properties. We show that subsequences of D-algebraic sequences, indexed by arithmetic progressions, satisfy ADEs of the same orders as the original sequences. Additionally, we discuss the special difference-algebraic nature of holonomic and $C^{2}$-finite sequences.
@@
Affine D-algebraic sequence, rationalizing difference polynomial, subsequence
@@ 1 Introduction
to accelerate the rate of convergence to $\lim_{n\to\infty}s(n)<\infty$. As discussed in [Wen89], this transformation has numerous applications in theoretical physics and other sciences [BJ75, Bre85]. This paper presents a general study of a class of sequences that is stable under such transformations. These sequences satisfy nonlinear rational or polynomial recursions, similar to the one presented in (1). Although we do not aim to describe an algorithm for numerical computations, some of our examples are motivated by sequence acceleration in numerical analysis.
@@
To give a glimpse of what our symbolic computations provide for Aitken’s $\Delta^{2}$ process, consider approximating $\sqrt{\ell},\,\ell>0,$ using the Babylonian method. The iteration is governed by the recursion
$s(n+1)=\frac{1}{2}\left(s(n)+\frac{\ell}{s(n)}\right),$ (2)
@@
with the chosen initial term $s(0)=s_{0}$. Using our result, we can systematically show that the Aitken transformation yields the sequence with initial term $t(0)=\frac{\ell(\ell+3s_{0}^{2})}{s_{0}(s_{0}^{2}+3\ell)}$ and recursion
$t(n+1)=\frac{2\,\ell\,t(n)}{t(n)^{2}+\ell}=\frac{2\,\ell}{t(n)+\frac{\ell}{t(n)}}.$ (3)
@@
1.1. Context. Following J. F. Ritt’s algebraic approach to differential equations [Rit50], the study of associated objects such as differential polynomials and differential varieties was further developed by Raudenbush [Rau34], Rosenfeld [Ros59], and, more comprehensively with modern mathematics, by Kolchin [Kol73]. For difference equations, which relate to the classical $\Delta$ operator defined as $\Delta f(x)\coloneqq f(x+1)-f(x)$, R. Cohn developed much of the theory [Coh65].
@@
A key result of R. Cohn states that every nontrivial ordinary algebraically irreducible difference polynomial admits an abstract solution [Coh48]. This theorem has important implications for sequences over an algebraically closed field, say $\mathbb{K}$, of characteristic zero. Indeed, specialized to the difference ring $\mathbb{K}^{\mathbb{N}}$ (with the monoid of natural integers $\mathbb{N}=\{0,1,\ldots\}$), Cohn’s theorem ensures the existence of sequences that are zeros of ordinary difference polynomial. We call them difference-algebraic (or D-algebraic) sequences and denote their set by $\mathbb{K}^{\aleph_{0}}$ (see Definition˜2.13).
The scarcity of practical algorithms for arithmetic computations with D-algebraic sequences can largely be attributed to the presence of zero-divisors in sequence rings. It is only recently that a theory for the zeros of algebraic (or polynomial) difference equations (ADEs) within these rings has been proposed [OPS20]. This is a non-trivial undertaking, as many established results from difference and differential fields do not readily translate to sequence rings [HP07]. Wibmer introduced a notion of dimension for systems of ADEs with solutions in sequence rings, showing that this dimension need not be an integer [Wib21]. Decidability results for solving systems of difference equations in sequences are discussed in [PSW20].
1.2. Related work. In the spirit of [Kau07], we employ difference algebra as a convenient language to investigate the closure properties of D-algebraic sequences. We aim to provide a computational framework for D-algebraic sequences from which known answers related to D-algebraic functions can also be addressed.
@@
D-algebraic functions [AEMSTT24, TT25] generalize functions defined by polynomial differential equations, encompassing classes such as D-finite functions [Kau23] and differentially definable functions [JPPS20] (see also [TTK21] for related power series). Similarly, D-algebraic sequences generalize sequences defined by polynomial equations involving their shifts and the index variable. These sequences arise naturally in computer science, for example, in the context of cost-register automata [ADD+13] and polynomial automata [BDSW17] over a unary alphabet. One also finds D-algebraic sequences in modeling, as illustrated with the discrete May-Leonard model by Roeger and Allen [RA04].
A particularly important subclass of D-algebraic sequences consists of rational recursive sequences, which are the zeros of difference polynomials linear in their highest shifts (or order terms) [CDBMP23, Ste04]. We refer to these polynomials as rationalizing difference polynomials.
Of course, for a given sequence $(s(n))_{n\in\mathbb{N}}$, the relation
$s(n+r)\,\mathsf{D}\!\left(s(n),s(n+1),\ldots,s(n+r_{1})\right)=\mathsf{N}\!\left(s(n),s(n+1),\ldots,s(n+r_{2})\right),$ (4)
with polynomials $\mathsf{N},\mathsf{D}$, and integers $r>r_{1},r_{2}$, is “meaningful” only if the sequence $(s(n))_{n\in\mathbb{N}}$ is not a zero of both the difference polynomials encoded by $\mathsf{N}$ and $\mathsf{D}$ simultaneously. This, however, does not affect the definition of a rationalizing difference polynomial; instead, it emphasizes the type of sequence solution. In this paper, we discuss a special type of so-called generic (sequence) solutions [Lev08, P. 32].
1.3. Contributions. We propose a formal definition for D-algebraic sequences. In the case of irreducible difference polynomials, this definition corresponds to special germs of their generic solutions. We then generalize this perspective to accommodate D-algebraic sequences with zeros of arbitrary difference polynomials.
It was shown in [TTW24] that any D-finite (holonomic or P-recursive) equation can be converted into a rationalizing ADE of order bounded by the sum of its holonomic degree and order. It turns out that the corresponding holonomic sequences are often generic zeros of the resulting ADE. We provide a corrected proof of this statement here, as an error was inadvertently introduced in the final version of [TTW24].
Another notable subclass of D-algebraic sequences is that of $C^{2}$-finite sequences, introduced by Jiminez-Pastor, Nuspl, and Pillwein in [JPNP23]. We establish a result analogous to Theorem 2 from [TTW24] to give a constructive proof of their D-algebraicity. This proof is based on a new algorithm, Algorithm˜2, which converts any $C^{2}$-finite equation into a rationalizing difference polynomial. Notably, our algorithm makes no explicit use of Gröbner bases or linear system solving. While we observe similarities with the resultant method from the differential case [JPPS20], the primary concern of that approach is not to derive an equation in which the highest order term appears linearly.
Note that the differential analogs of these conversions are far more straightforward. Indeed, in the differential case, the algorithm for D-finite functions essentially involves eliminating the independent variable via the resultant method, followed by computing the first derivative of the resulting differential polynomial. Similarly, unlike $D^{2}$-finite equations, the conversion of $C^{2}$-finite equations into rational recursions appears to require different techniques.
Regarding arithmetic operations, while the computational theory for the differential case is relatively well-established (see, e.g., [BLOP95, Rob14, vDH19, TT25] and references therein), an effective theory for the arithmetic of (nonlinear) difference-algebraic sequences has received less attention. This may also explain the lack of results concerning D-algebraic subsequences, and thus motivates the present work.
This paper demonstrates that D-algebraic sequences behave well under field operations, provided that mild assumptions are made about their consecutive terms. For illustration purposes, consider the following algebraic difference equation satisfied by $(n!)_{n\in\mathbb{N}}$:
$s\!\left(n+2\right)=\frac{s\!\left(n+1\right)\left(s\!\left(n\right)+s\!\left(n+1\right)\right)}{s\!\left(n\right)}.$ (5)
This equation may be obtained using the algorithm from [TTW24] with the minimal D-finite equation satisfied by $n!$ as input. We may use this recursion to construct an ADE satisfied by $u(n)\coloneqq\frac{n!}{n!+1}$, which is not D-finite. For $s(n)$ satisfying (5) with $s(0)=s(1)=1$, we have that $u(n)=~\frac{s(n)}{s(n)+1}$, a rational expression in $s(n)$. Provided that
$s(n)s(n+1)s(n+2)(s(n)+1)(s(n+1)+1)(s(n+2)+1)\neq 0,$ (6)
Algorithm˜1 computes the following ADE for $u(n)$, which we write as a recursion to make the coefficient of the order term explicit:
$u\!\left(n+2\right)=\frac{u\!\left(n+1\right)\left(2u\!\left(n+1\right)u\!\left(n\right)-u\!\left(n+1\right)-u\!\left(n\right)\right)}{u\!\left(n+1\right)^{2}u\!\left(n\right)-u\!\left(n+1\right)^{2}+u\!\left(n+1\right)u\!\left(n\right)-u\!\left(n\right)}.$ (7)
By considering generic zeros of difference polynomials, we neglect all sequence solutions of the ADE in (7) that vanish the denominator on its right-hand side. The conditions made in (6) arise from shifts of the denominators in (5) and the relation between $s(n)$ and $u(n)$, which are used during the Gröbner bases computations. Again, this condition may be simplified for generic sequences satisfying (5). However, these conditions can be applied to practical computations involving actual sequences. For a specific $n_{0}\in\mathbb{N}$, it is required to verify that all these conditions are satisfied before using (7) to compute the subsequent terms of $(u(n))_{n\geq n_{0}}$. Since for $(n!)_{n\in\mathbb{N}}$, the non vanishing of the denominator in (7) (for $u(n)$) and the conditions in (6) are always satisfied, all terms of $(u(n))_{n\in\mathbb{N}}$ may be computed with (7) and its first two initial values. In essence, our algorithmic computations with D-algebraic sequences rely on these kinds of generic assumptions, which are highly dependent on the initial terms of these sequences. The latter already relates to a longstanding decision problem for $C$-finite sequences [OW12]. However, we will not discuss this aspect further, as our interest in this paper lies only in the symbolic constructions with difference polynomials.
In addition to field operations and completing the work in [TTW24], we prove that D-algebraic sequences are closed under partial sums, partial products, radicals (one may assume complex values and appropriate choices for branches), and composition with arithmetic progressions. This last result can be extended to composition with broader classes of strictly increasing integer sequences. We are unaware of comparable results for such compositions, and we consider our method for computing equations for D-algebraic subsequences another key contribution.
Our method for all these operations is based on the decomposition-elimination-prolongation method developed by Ovchinnikov, Pogudin, and Vo [OPV22], which was adapted to the difference algebra setting in [OPS20] (see also [GVDHYZ09]). We construct difference polynomials for D-algebraic closure properties by adapting the result from [TT25], which proposes a method to deal with difference polynomials that are nonlinear in their highest shifts.
@@ 2 Definitions
We primarily consider the $\sigma$-ring $(\mathbb{K},\sigma)$, where $\mathbb{K}$ is an algebraically closed field of characteristic zero and $\sigma$ an endomorphism in $\mathbb{K}$.
@@
[…]
@@
The order of a difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, denoted $\operatorname{ord}(p)$, is the maximum $j\in\mathbb{N}$ such that $\sigma^{j}(x)$ appears in $p$. The degree of $p$ denoted $\deg(p)$ is the total degree of $p$ as a polynomial in the algebraic polynomial ring $\mathbb{K}[x,\sigma(x),\ldots,\sigma^{\operatorname{ord}(p)}(x)]$. We say that $p$ is rationalizing when $p$ is linear in its highest shift term, i.e., the degree of $p$ with respect to $\sigma^{\operatorname{ord}(p)}(x)$ is $1$.
To any difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, we associate a recurrence equation called algebraic difference equation obtained by the equality $p=0$. For a rationalizing difference polynomial, the associated recurrence equation is equivalent to a rational recursion.
@@
The difference polynomial associated to (5) is $\sigma^{2}(x)x-\sigma(x)(x+\sigma(x)),$ of order $2$ and degree $2$. For (7), the corresponding difference polynomial is of order $2$ and degree $4$. $\blacksquare$
In classical algebraic geometry, zeros of polynomials over $\mathbb{K}$ correspond to points in $\mathbb{K}^{N}$, the affine $N$-space over $\mathbb{K}$, for some fixed $N\in\mathbb{N}\setminus\{0\}$, defining algebraic sets or varieties with the ideals of those polynomials. In our setting, the corresponding points may be regarded with denumerable coordinates as we want any zero of $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ to have coordinates that vanish $\sigma^{j}(p)$, for all $j\in\mathbb{N}$. Every such point defines a sequence over $\mathbb{K}$. We denote the set of such sequences over $\mathbb{K}$ by $\mathbb{K}^{\aleph_{0}}$, where $\aleph_{0}$ is aleph zero, the cardinality of $\mathbb{N}$. This makes sense of the passage from a finite to an infinite (countable) number of coordinates. Note that $\mathbb{K}^{\aleph_{0}}\subset\mathbb{K}^{\mathbb{N}}$, and the inclusion is strict since $(n^{n})_{n\in\mathbb{N}}$ is not a zero of a difference polynomial. The operations in $\mathbb{K}^{\aleph_{0}}$ are inherited from the difference ring $\mathbb{K}^{\mathbb{N}}$ defined with the shift endomorphism.
@@
Let $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ and $(s(n))_{n\in\mathbb{N}}\in(\mathbb{K}^{\aleph_{0}},\sigma^{\prime})$ as previously defined. We say that $(s(n))_{n\in\mathbb{N}}$ is a zero of $p$ if, under the unique homomorphism of difference rings $\mathcal{D}_{\sigma}(\mathbb{K},x)\longrightarrow~(\mathbb{K}^{\aleph_{0}},\sigma^{\prime})$ given by the extension that sends $x$ to $(s(n))_{n\in\mathbb{N}}$, $p$ is sent to 0. We interpret the fact that $(s(n))_{n\in\mathbb{N}}$ is a zero of $p$ by “$p(s(n))=0$ for all non-negative integers nn” or $p((s(n))_{n\in\mathbb{N}})=0$” (or simply $p(s(n))=0$ with unspecified $n$).
Several sequences are zeros of difference polynomials, but in some cases, being a zero of a difference polynomial does not particularly add value to the knowledge of the sequence. According to Definition˜2.6, even the sequence of Bernoulli numbers satisfies an algebraic difference equation. Indeed, consider the difference polynomial
$p\coloneqq 5\,\sigma^{3}(x)\,x-6\,\sigma^{2}(x)\sigma(x)+\sigma(x)\,x.$ (8)
Recall that the $n$th Bernoulli number $B_{n}$ may be computed by the formulas:
$\begin{split}&B_{0}=1,\,B_{2n}=\frac{(-1)^{n+1}2\,(2n)!}{(2\pi)^{2n}}\zeta(2n),\,n\geq 1,\\ &B_{1}=-\frac{1}{2},\,B_{2n+1}=0,\,n\geq 1,\end{split}$ (9)
where $\zeta$ is the Riemann Zeta function. One verifies that for all $n\in\mathbb{N}$, $p(B_{n})=0$. However, the only terms of $(B_{n})_{n\in\mathbb{N}}$ that can be deduced from $p$ are the terms $B_{2n},n\leq 2$ and $B_{2n+1},n\geq 0$. When such a phenomenon happens for a given sequence and a difference polynomial, we say that the sequence is not a generic zero of that difference polynomial. Let us make this statement more precise.
Definition 2.7.
Let $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$. The initial $\mathcal{I}_{p}$ of $p$ is the leading coefficient of $p$ viewed as a univariate polynomial in $\sigma^{\operatorname{ord}(p)}(x)$.
Proposition 2.8.
Let $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)\setminus\mathbb{K}$. We have $\operatorname{ord}(\mathcal{I}_{p})<\operatorname{ord}(p)$ and $\deg(\mathcal{I}_{p})<\deg(p)$.
Proof.
The inequality for the order is immediate from the definition. For the degree, let $r=\operatorname{ord}(p)$ and $m\coloneqq\deg_{\sigma^{r}(x)}(p)>0$ be the degree of $p$ in $\sigma^{r}(x)$. From the definition of $\mathcal{I}_{p}$, it follows that there exists $p_{1}\in\mathcal{D}_{\sigma}(\mathbb{K},x),\deg_{\sigma^{r}(x)}(p_{1})<m,\operatorname{ord}(p_{1})\leq r$ such that
$p=\mathcal{I}_{p}\left(\sigma^{\operatorname{ord}(p)}(x)\right)^{m}+p_{1},$ (10)
showing that $\deg(p)\geq\deg(\mathcal{I}_{p})+m>\deg(\mathcal{I}_{p})$. ∎
Definition 2.9.
Let $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$. A sequence $(s(n))_{n\in\mathbb{N}}$ is a generic zero of $p$ if it is a zero of all $P$ in the difference ideal <p><p> and not a zero of any $P\notin<p>$.
Proposition 2.10.
If a sequence $(s(n))_{n\in\mathbb{N}}$ is a generic zero of an algebraically irreducible difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, then $\mathcal{I}_{p}((s(n))_{n\in\mathbb{N}})\neq 0$.
Proof.
If $\mathcal{I}_{p}((s(n))_{n\in\mathbb{N}})=0$ then $\mathcal{I}_{p}\in<p>$. Since $p$ is irreducible, there exist difference polynomials $q_{0},q_{1},\ldots,q_{k}\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ such that
$\mathcal{I}_{p}=\sum_{j=0}^{k}q_{j}\sigma^{j}(p).$ (11)
Using an orderly ranking on $R\coloneqq\mathbb{K}[x,\sigma(x),\ldots,\sigma^{k+\operatorname{ord}(p)}(x)]$, the coefficients $q_{j}$’s can be interpreted as the quotients in a reduction (division) of $\mathcal{I}_{p}$ with respect to the triangular set $\{\sigma^{j}(p),j=0\ldots,k\}$ generating the truncation of <p><p> in $R$. By the reduction algorithm, we must have that
$\deg(\mathcal{I}_{p})=\deg(q_{k}\,\sigma^{k}(p))=\deg(q_{k})+\deg(p),$
using the correspondence condition between the leading terms on both sides in (11).
However, by Proposition˜2.8 we know that $\deg(\mathcal{I}_{p})<\deg(p)$. Therefore we must have $\deg(q_{k})<~0$, thus $q_{k}=0$. Repeating this reasoning with the remaining $q_{j}$’s, $0\leq j<k$, leads to $\mathcal{I}_{p}=0$, a contradiction since $p$ is irreducible. ∎
The following theorem may be regarded as a corollary of R. Cohn’s existence theorem for ordinary difference polynomials.
@@
[Coh48, Theorem IV] Every irreducible difference polynomial in $\mathcal{D}_{\sigma}(\mathbb{K},x)\setminus\mathbb{K}$ has a generic zero.
Following Cohn’s original proof, the core idea of a generic zero—as illustrated by Proposition˜2.10 and Theorem˜2.11—is that all subsequent terms of such a sequence can be computed from a corresponding difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ and its first $\operatorname{ord}(p)$ initial terms. This property, however, is the essential problem that hampers the definition of the Bernoulli sequence using the irreducible polynomial $p$ from (8).
The classical definition of a generic zero does not fully capture this idea of computability. We resolve this by introducing the class of D-algebraic sequences. To simplify the understanding of our main definition, we first consider the case of irreducible difference polynomials.
Definition 2.12 (D-algebraicity with irreducible difference polynomial).
Let $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ be irreducible. A sequence $(s(n))_{n\in\mathbb{N}}$ is difference-algebraic (or D-algebraic) over $\mathbb{K}$ with defining polynomial $p$, if $\mathcal{I}_{p}(s(n))\neq 0$ for sufficiently large $n$.
Whether the Bernoulli sequence is D-algebraic or not is not proven in this paper. However, from Definition˜2.12 we know that the difference polynomial $p$ in (8) cannot be used to define the Bernoulli sequence since $\mathcal{I}_{p}(B_{2n+1})=0,n\geq 1$. In other words, for an irreducible difference polynomial $p$ to define a D-algebraic sequence $(s(n))_{n\in\mathbb{N}}$, we must have $|\{n\in\mathbb{N},\mathcal{I}_{p}(s(n))=0\}|<\infty.$ We refer to this condition as the regularity condition.
In general, to show that a given sequence is D-algebraic, it is enough to show that all its terms can be obtained by finding the roots of some univariate polynomials resulting from evaluating a difference polynomial with its previous terms. Generalizing Definition˜2.12 for arbitrary difference polynomials, we obtain the definition below.